given the following information U(x y)=x^0.5,y^0.5,the initial price of good X is $10 and good Y is $5:consumers income $1000.calculate the optimal bundle at initial bundle with good X on horizontal axis
1
Expert's answer
2020-06-16T12:15:21-0400

Optimal bundle will occur where:



"\dfrac{MU_x}{MU_y} = \dfrac{P_x}{P_y}"

The utility function is:



"U(x, y)=x^{0.5}y^{0.5}"

Therefore:



"MU_x = 0.5x^{-0.5}y^{0.5}\\[0.3cm]
MU_y = 0.5x^{0.5}y^{-0.5}"

In our question, "P_x = \$10, \quad P_y = \$5" . Therefore:



"\dfrac{0.5x^{-0.5}y^{0.5}}{0.5x^{0.5}y^{-0.5}} = \dfrac{10}{5}\\[0.3cm]
\dfrac{y}{x} = 2\\[0.3cm]
y = 2x......\text{Eqn 1}"

The consumer's income is $1,000. Therefore, the budget constraint is:



"1000 = 10x + 5y"

Substituting equation 1 into the budget constraint:



"1000 = 10x + 5(2x)\\[0.3cm]
1000 = 20x\\[0.3cm]
\color{red}{x^* = 50}"

Since "y = 2x" , then:



"y^* = 2(50)\\[0.3cm]
\color{red}{y^* = 100}"